# Aethorix v1.0: An Integrated Scientific AI Agent for Scalable Inorganic Materials Innovation and Industrial Implementation

### v2

@@ Abstract
Artificial Intelligence (AI) is redefining the frontiers of scientific domains, ranging from drug discovery[1, 2, 3, 4] to meteorological modeling[5, 6, 7], yet its integration within industrial manufacturing remains nascent and fraught with operational challenges. To bridge this gap, we introduce Aethorix v1.0, an AI agent framework designed to overcome key industrial bottlenecks, demonstrating state-of-the-art performance in materials design innovation and process parameter optimization. Our tool is built upon three pillars: a scientific corpus reasoning engine that streamlines knowledge retrieval and validation, a diffusion-based generative model for zero-shot inverse design, and specialized interatomic potentials that enable faster screening with abab initioinitio fidelity. We demonstrate Aethorix’s utility through a real-world cement production case study, confirming its capacity for integration into industrial workflows and its role in revolutionizing the design-make-test-analyze loop while ensuring rigorous manufacturing standards are met.
@@ 1 Introduction
The Fourth Industrial Revolution (Industry 4.0) has ushered in a new era of manufacturing defined by the pervasive integration of cyber-physical systems, artificial intelligence (AI), and digital fabrication.[8, 9] This paradigm enables mass personalization, smart products, and data-driven agility,[10] yet it has also exposed a critical asymmetry: while digital capabilities advance exponentially, the backbone of (bio)materials innovation remains entrenched in conventional, human-intuition-driven methodologies.[11] This bottleneck is acutely highlighted by the vastness of chemical space; the $10^{60}$ drug-like compounds, for instance, exceed human cognitive capacity by tens of orders of magnitude.[12] While the development of self-driving laboratories represents a significant step forward,[13, 14, 11, 15] their output often struggles to translate into realistic, manufacturable products, due in part to an overreliance on idealized computational data that fails to capture real-world non-ideality.[16]
@@
Interviews with industrial stakeholders across the United States and China reveal that the core challenge lies not only in materials innovation but, more critically, in the manufacturing phase, which are governed by intricate, multi-step, multi-scale thermophysical and chemokinetic phenomena.[17, 18] Here, high-fidelity atomistic simulations are computationally prohibitive to deploy at the spatial and temporal scales of continuous production, whereas agnostic data-driven methodologies lack embedded physicochemical realism, rendering them unreliable for predictive control and scale-up due to their failure to generalize beyond calibrated regimes. Bridging this gap requires an integrated framework that synergistically combines operational data with physics-based models.[19]
@@
However, the current deployment of AI in scientific domains remains largely confined to auxiliary tasks such as literature mining, information retrieval, and decontextualized property analysis.[20, 21, 22, 23, 24, 25, 26, 27] Achieving higher-order scientific reasoning and innovation is hampered by several core limitations: (1) data fragmentation across material classes impedes cross-domain generalization;[28] (2) the high-dimensional, context-dependent nature of material behavior, governed by multivariate conditions such as temperature, pressure, and local chemical environment, renders decontextualized predictions scientifically incomplete;[29] and (3) the multi-scale character of physicochemical phenomena, spanning femtosecond electronic excitations to centennial geological processes, presents a major obstacle for unified modeling.[30]
@@
An emerging frontier is the transition from passive computational predictive models to active, goal-directed inverse design.[31, 32, 33, 34] This trending methodology would enable the autonomous navigation of chemical space toward desired properties, moving beyond the brute-force, trial-and-error approach of high-throughput screening, which merely accelerates evaluation rather than intelligently guiding exploration.[32] Overcoming these intertwined bottlenecks in both materials innovation and process optimization is critical for vertical industries to meet accelerating market demands and sustain a competitive advantage.
@@ 4.1 First-Principles Atomistic Data
The foundational atomistic dataset is constructed from high-fidelity, first-principles calculations based on the density functional theory (DFT). These are the computational ground truth for mapping structure-energy relationships, which are fundamental to the generative and predictive modules within the agent workflow. Our work leverages and extends large-scale, publicly available benchmarks. This includes the Open Molecule 2025 (OMol 25) [35] (100M+ single-point calculations), the Open Materials 2024 (OMat24) [36] (110M+ single-point calculations), and the Open Molecular Crystals 2025 (OMC25) [37] (27M+ single-point calculations) datasets as our primary bulk structure repositories. For surface-specific phenomena, we utilize the Open Catalyst 2020 (OC20) [38] ($\sim$265M single-point calculations) and the Open Catalyst 2022 (OC22) [39] ($\sim$9.8M single-point calculations) datasets. All datasets consist of an ensemble of unrelaxed and relaxed structures, spanning a wide range of atom types, along with their corresponding DFT-calculated total energies, atomic forces, and stress tensors across diverse adsorbate species. Additional datasets are supplied via our Structure Generator and in-house DFT calculations for specific systems of industrial interest, fulfilling the undersampling gaps for targeted compositions and configurations within public repositories. To ensure data consistency, the exchange-correlation functionals and pseudopotentials used for these supplemental calculations are meticulously aligned with those of the referenced public datasets. All in-house calculations are performed with the Quantum Espresso software package[40, 41] and the Vienna abab initioinitio simulation package (VASP)[42].
@@ 5 Scientific Literature Review
The scientific literature review is implemented via a Retrieval-Augmented Generation (RAG) pipeline architected around a Large Language Model (LLM). This module ingests pre-processed user queries to conduct multi-stage, vector-based retrieval from a curated scientific corpus (enabled by SciSpace[43]). The retrieval subsystem utilizes open-source, domain-specific text representation models to execute parallel dense searches against a low-latency vector database. This is powered by SPECTER[44] for document-level embeddings and ColBERT[45] for late-interaction, contextualized token embeddings, which collectively yielding an initial candidate set of several hundred documents. The candidate set is subsequently refined by a neural re-ranker that optimizes for relevance using a composite of features, narrowing the results to the top-k most pertinent passages. The LLM then synthesizes these retrieved, citation-grounded contexts to generate a coherent literature review, thereby constraining generative hallucinations and ensuring the output’s verifiability against the source corpus.
@@ 6 Structure Generation
All atomistic structural modeling is managed by the Atomic Simulation Environment (ASE) [46]. The source of initial structures is contingent upon the design objective: for targeted exploration of known crystal systems, initial configurations are sourced from the Materials Project[47] database with rational compositional and configurational modifications, whereas for dede novonovo design, the goal is accomplished using a symmetry-aware, diffusion-based generative model. The implementation is adapted from MatterGen[32], which is architected for zero-shot generation of novel and thermodynamically stable inorganic materials. In this scheme, periodic crystal materials are represented by $M=(A,X,L)$, where $A=(a^{1},a^{2},...,a^{n})^{T}\in\mathbb{A}^{n}$ represents the atom species inside the unit cell, $X=(x^{1},x^{2},...,x^{n})\in[0,1)^{3\times n}$ denotes the fractional coordinates of atoms in the unit cell, and $L=(l^{1},l^{2},l^{3})\in\mathbb{R}^{3\times 3}$ is the lattice with three vector components. During the forward diffusion (noising) process, $A$, $X$, and $L$ are independently corrupted toward a physically meaningful prior distribution as follows:
@@
To enhance the efficiency of the downstream high-throughput screening process, a property-guided generation capability is integrated. This is achieved by fine-tuning the unconditional score network by augmenting its node representations with property-specific features through trainable adapter modules.[48] The property-augmented node hidden representation is given by:
@@ 7 Structure Optimization
DFT is a standard tool for the ground-state energetics calculations of crystal structures, which can be followed by geometric optimization to determine the equilibrium atomic configuration. This is typically achieved using the limited memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS)[49] or the fast inertial relaxation engine (FIRE)[50] algorithm. However, its computational cost makes it difficult to scale for high-throughput industrial applications, necessitating the development of surrogate models that reproduce abab initioinitio accuracy at orders-of-magnitude lower computational expense. Machine-learned interatomic potentials (MLIPs) utilize parametric functionals to represent interatomic interactions. Trained on atomic structure data, MLIPs are developed through a graph neural network (GNN)-based architecture, which directly maps atomic configurations to their corresponding potential energy surface (PES), without the need for handcrafted feature descriptors. Many state-of-the-art (SOTA) GNNs, such as MACE[51], NequIP[52], eSEN[53], and PaiNN[54], have been developed to leverage equivariant message passing to achieve high data efficiency. Our agent integrates the open-source Universal Models of Atoms–Medium (UMA-M)[55], which comprises five specialized models, each uniquely designed for distinct categories of material systems. While the models have exhibited remarkable generalizability across a large chemical space, prediction accuracy for novel compositions may still be constrained by the inherent limitations of the training data, which lack a proper description of charge, spin, and strong $d$-orbital correlations. For specialized systems where these effects are dominant, our agent develops specialized MLIPs, which are typically trained on only a few atom types, but they can be rapidly fine-tuned and uncertainty quantified via ensemble variance methods. Training of MLIPs is implemented in a multi-target learning scheme, where the node features are optimized to simultaneously predict multiple properties. The loss function is given by:
@@
where $\alpha_{E}$, $\alpha_{F}$, and $\alpha_{S}$ are the weight coefficients for total energy, atomic force, and virial stress, respectively, $N$ denotes the number of training samples, $X_{i}$ and $\hat{X}_{i}$ ($X=E,\mathbf{F},\sigma$) are the true and predicted targets for structure $i$, respectively. Optimized hyperparameters are often determined using Bayesian optimization methods.[56] MLIP ensembles are constructed by training multiple models on identical datasets but with different parameter initializations. The variance in predictions across the ensemble provides a rigorous, $a$ prioripriori metric for assessing model confidence and identifying extrapolation beyond the trained domain[57]. Active learning is implemented to ensure transferability of specialized models across the desired chemical space.[58, 59, 60] The training dataset is iteratively enriched by incorporating new structures that are generated from the conditioned atom types and have the highest epistemic uncertainty.
@@ 8 Physics-Informed Prediction
Many emergent macroscale properties of materials cannot be simply derived from a single DFT geometric optimization calculation, but instead require multiple DFT calculations integrated with physics-informed analysis. Hence, the computational cost can grow exponentially with increasing system size and project scope. In contrast, MLIPs can unlock the unprecedented efficiency in predictive screening while preserving near-$ab$ initioinitio fidelity. Figure 2 contrasts the computational cost scaling between DFT and MLIPs across a range of system sizes for various simulation tasks. These performance benchmarks are averaged from the estimated calculation convergence time observed across a series of Mg-Si-O-H chemical systems. The computational cost of obtaining emergent properties via MLIP inference is 103 $\sim$ 105 times lower than equivalent DFT calculations. It should be noted that the exact magnitude of this advantage can vary with the chemical system, the resolution of the computational methods employed ($e.g.,$ the number of intermediate images in nudged elastic band calculations), and the simulation timescale ($e.g.,$ for molecular dynamics trajectories), yet it consistently reflects a significant reduction in computational expense for MLIP. Crucially, this cost reduction renders many prohibitively expensive calculations ($e.g.,$ phonon vibrations) feasible for high-throughput screening.
@@ 9.1 Problem Framing
Herein, we present a use case of Aethorix v1.0 in resolving a persistent quality control challenge in Portland cement manufacturing. As detailed in Figure 3, the objective of the example use case is to systematically reduce the day-to-day production variability in the compressive strength across both early-age (1-3 day) and long-term (28 day) curing regimes. After consolidating the predominant expert perspectives from a corpus of 18 scholarly works published within the last five years,[61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78] the integrated SciSpace module summarized leading causal factors and proposed established mitigation strategies for raw materials engineering and process control optimization (Figure 3a). Following user interaction, our agent formulated actionable plans for solution development (Figure 3b), prompting the user to select an implementation path based on their prioritized needs. It is noteworthy that the framework supported further interactive dialogue to elaborate on the underlying design principles of each proposed solution.
@@ 9.2 Phase Stabilization through Atomistic Engineering
The elevated calcination temperature of 1730 K necessitated explicit consideration of entropic contributions to phase stability. C3S and C2S are the dominant compositions in cement clinkers. The computed Gibbs free energies of formation for polymorphic C3S and C2S phases at finite temperatures are presented in Figure 4e–g and Figure S1, respectively. The thermodynamic landscape reveals a broad distribution of triclinic and orthorhombic symmetries across both systems. The distribution shifted significantly between room temperature and 1730 K. While experimental identification of exact C3S and C2S crystallographic symmetries remains challenging, our predicted most stable phases exhibit good consistency with the reported industrial clinker polymorphs.[79] At 1730 K, the equilibrium lattice parameters of the most thermodynamically stable C3S and C2S polymorphs are shown as follows:
@@
The early-age properties and long-term performance of concrete are directly governed by the hydration behavior of its clinkers.[80, 81] However, the hydration of clinkers involves complex, coupled reaction mechanisms ($e.g.,$ ionic dissolution, degree of hydroxylation, evolution of surface structure, and chemical bond formation),[81, 82] making the modeling of the full process at atomistic levels prohibitively expensive. Our agent captured the bulk ionic mobility to approximate hydration reactivity of clinkers. Figure 4f shows the ionic self-diffusivity distribution for the thermodynamically most stable C3S polymorphs, quenched from 1730 K. The self-diffusivity values were derived from short-time molecular dynamics simulations (1 ps) at 300 K via the Einstein relation:
@@
where $N$ is the number of atoms, $d$ is the dimensionality of diffusion, $\mathbf{r}_{i}$ is the position of atom $i$ at $t$, and the angular brackets denote an ensemble or time average. The self-diffusivity was found to vary by 1–2 orders of magnitude across all thermodynamically most stable polymorphs. Those with low ionic mobility ($e.g.,$ ID 01) were hence identified undesirable and their production was targeted for suppression during synthesis. Figure 4g and S1c show the phase stabilization effect due to presence of low concentration ($\sim$10 mol%) of dopant ions in clinker crystal lattices at 1730 K. The results revealed varying dopant site occupancy preferences across polymorphs, which deviate from empirical trends.[83, 84, 85] For instance, Na preferentially substitutes the Ca site in the C2S ID10 lattice, whereas it favors the Si site in the C3S ID25 lattice. Most of these substitutions resulted in a reduced free energy at 1730 K, Notable exceptions, such as S or P substituting for Ca, are likely due to the significant charge mismatch introduced by the dopant, which destabilizes the lattice. These are in agreement with experimental observations.[79] Crucially, K doping provided the strongest stabilization of all dopants for the desired polymorphs, while having a negligible effect on the undesired ones. It was further verified that incorporation of K does not impede ion mobility within the desired phases. Therefore, it was proposed that industrial clinker production strategically maintain K impurity levels at 5–10 mol% to favor the formation of hydration-reactive polymorphs, a direct route to improving the performance of concrete from early ages to long-term service.

### v3

@@ Abstract
Artificial Intelligence (AI) is redefining the frontiers of scientific domains, ranging from drug discovery[jumper2021highly, blanco2023role, zhang2025artificial, jayatunga2022ai] to meteorological modeling[dewitte2021artificial, conti2024artificial, bi2023accurate], yet its integration within industrial manufacturing remains nascent and fraught with operational challenges. To bridge this gap, we introduce Aethorix v1.0, an AI agent framework designed to overcome key industrial bottlenecks, demonstrating state-of-the-art performance in materials design innovation and process parameter optimization. Our tool is built upon three pillars: a scientific corpus reasoning engine that streamlines knowledge retrieval and validation, a diffusion-based generative model for zero-shot inverse design, and specialized interatomic potentials that enable faster screening with abab initioinitio fidelity. We demonstrate Aethorix’s utility through a real-world cement production case study, confirming its capacity for integration into industrial workflows and its role in revolutionizing the design-make-test-analyze loop while ensuring rigorous manufacturing standards are met.
@@ 1 Introduction
The Fourth Industrial Revolution (Industry 4.0) has ushered in a new era of manufacturing defined by the pervasive integration of cyber-physical systems, artificial intelligence (AI), and digital fabrication.[petrillo2018fourth, philbeck2018fourth] This paradigm enables mass personalization, smart products, and data-driven agility,[schwab2024fourth] yet it has also exposed a critical asymmetry: while digital capabilities advance exponentially, the backbone of (bio)materials innovation remains entrenched in conventional, human-intuition-driven methodologies.[pyzer2022accelerating] This bottleneck is acutely highlighted by the vastness of chemical space; the $10^{60}$ drug-like compounds, for instance, exceed human cognitive capacity by tens of orders of magnitude.[reymond2012enumeration] While the development of self-driving laboratories represents a significant step forward,[szymanski2023autonomous, dai2024autonomous, pyzer2022accelerating, seifrid2022autonomous] their output often struggles to translate into realistic, manufacturable products, due in part to an overreliance on idealized computational data that fails to capture real-world non-ideality.[peplow2025ai]
@@
Interviews with industrial stakeholders across the United States and China reveal that the core challenge lies not only in materials innovation but, more critically, in the manufacturing phase, which are governed by intricate, multi-step, multi-scale thermophysical and chemokinetic phenomena.[zhao2020multiscale, oliveira2024learning] Here, high-fidelity atomistic simulations are computationally prohibitive to deploy at the spatial and temporal scales of continuous production, whereas agnostic data-driven methodologies lack embedded physicochemical realism, rendering them unreliable for predictive control and scale-up due to their failure to generalize beyond calibrated regimes. Bridging this gap requires an integrated framework that synergistically combines operational data with physics-based models.[blakseth2022combining]
@@
However, the current deployment of AI in scientific domains remains largely confined to auxiliary tasks such as literature mining, information retrieval, and decontextualized property analysis.[haristiani2019artificial, miret2024llms, kim2024large, cavanagh2024smileyllama, tang2025matterchat, xiao2024proteingpt, pal2023chatgpt, kon2025curie] Achieving higher-order scientific reasoning and innovation is hampered by several core limitations: (1) data fragmentation across material classes impedes cross-domain generalization;[schmidt2019recent] (2) the high-dimensional, context-dependent nature of material behavior, governed by multivariate conditions such as temperature, pressure, and local chemical environment, renders decontextualized predictions scientifically incomplete;[gurnani2024ai] and (3) the multi-scale character of physicochemical phenomena, spanning femtosecond electronic excitations to centennial geological processes, presents a major obstacle for unified modeling.[butler2018machine]
@@
An emerging frontier is the transition from passive computational predictive models to active, goal-directed inverse design.[wang2022inverse, zeni2025generative, fuhr2022deep, menon2022generative] This trending methodology would enable the autonomous navigation of chemical space toward desired properties, moving beyond the brute-force, trial-and-error approach of high-throughput screening, which merely accelerates evaluation rather than intelligently guiding exploration.[zeni2025generative] Overcoming these intertwined bottlenecks in both materials innovation and process optimization is critical for vertical industries to meet accelerating market demands and sustain a competitive advantage.
@@ 4.1 First-Principles Atomistic Data
The foundational atomistic dataset is constructed from high-fidelity, first-principles calculations based on the density functional theory (DFT). These are the computational ground truth for mapping structure-energy relationships, which are fundamental to the generative and predictive modules within the agent workflow. Our work leverages and extends large-scale, publicly available benchmarks. This includes the Open Molecule 2025 (OMol 25) [levine2025open] (100M+ single-point calculations), the Open Materials 2024 (OMat24) [barroso2024open] (110M+ single-point calculations), and the Open Molecular Crystals 2025 (OMC25) [gharakhanyan2025open] (27M+ single-point calculations) datasets as our primary bulk structure repositories. For surface-specific phenomena, we utilize the Open Catalyst 2020 (OC20) [chanussot2021open] ($\sim$265M single-point calculations) and the Open Catalyst 2022 (OC22) [tran2023open] ($\sim$9.8M single-point calculations) datasets. All datasets consist of an ensemble of unrelaxed and relaxed structures, spanning a wide range of atom types, along with their corresponding DFT-calculated total energies, atomic forces, and stress tensors across diverse adsorbate species. Additional datasets are supplied via our Structure Generator and in-house DFT calculations for specific systems of industrial interest, fulfilling the undersampling gaps for targeted compositions and configurations within public repositories. To ensure data consistency, the exchange-correlation functionals and pseudopotentials used for these supplemental calculations are meticulously aligned with those of the referenced public datasets. All in-house calculations are performed with the Quantum Espresso software package[giannozzi2009quantum, giannozzi2017advanced] and the Vienna abab initioinitio simulation package (VASP)[hafner2008ab].
@@ 5 Scientific Literature Review
The scientific literature review is implemented via a Retrieval-Augmented Generation (RAG) pipeline architected around a Large Language Model (LLM). This module ingests pre-processed user queries to conduct multi-stage, vector-based retrieval from a curated scientific corpus (enabled by SciSpace[jain2024scispace]). The retrieval subsystem utilizes open-source, domain-specific text representation models to execute parallel dense searches against a low-latency vector database. This is powered by SPECTER[cohan2020specter] for document-level embeddings and ColBERT[santhanam2021colbertv2] for late-interaction, contextualized token embeddings, which collectively yielding an initial candidate set of several hundred documents. The candidate set is subsequently refined by a neural re-ranker that optimizes for relevance using a composite of features, narrowing the results to the top-k most pertinent passages. The LLM then synthesizes these retrieved, citation-grounded contexts to generate a coherent literature review, thereby constraining generative hallucinations and ensuring the output’s verifiability against the source corpus.
@@ 6 Structure Generation
All atomistic structural modeling is managed by the Atomic Simulation Environment (ASE) [larsen2017atomic]. The source of initial structures is contingent upon the design objective: for targeted exploration of known crystal systems, initial configurations are sourced from the Materials Project[jain2013commentary] database with rational compositional and configurational modifications, whereas for dede novonovo design, the goal is accomplished using a symmetry-aware, diffusion-based generative model. The implementation is adapted from MatterGen[zeni2025generative], which is architected for zero-shot generation of novel and thermodynamically stable inorganic materials. In this scheme, periodic crystal materials are represented by $M=(A,X,L)$, where $A=(a^{1},a^{2},...,a^{n})^{T}\in\mathbb{A}^{n}$ represents the atom species inside the unit cell, $X=(x^{1},x^{2},...,x^{n})\in[0,1)^{3\times n}$ denotes the fractional coordinates of atoms in the unit cell, and $L=(l^{1},l^{2},l^{3})\in\mathbb{R}^{3\times 3}$ is the lattice with three vector components. During the forward diffusion (noising) process, AA, $X$, and $L$ are independently corrupted toward a physically meaningful prior distribution as follows:
@@
To enhance the efficiency of the downstream high-throughput screening process, a property-guided generation capability is integrated. This is achieved by fine-tuning the unconditional score network by augmenting its node representations with property-specific features through trainable adapter modules.[dunn2020benchmarking] The property-augmented node hidden representation is given by:
@@ 7 Structure Optimization
DFT is a standard tool for the ground-state energetics calculations of crystal structures, which can be followed by geometric optimization to determine the equilibrium atomic configuration. This is typically achieved using the limited memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS)[liu1989limited] or the fast inertial relaxation engine (FIRE)[bitzek2006structural] algorithm. However, its computational cost makes it difficult to scale for high-throughput industrial applications, necessitating the development of surrogate models that reproduce abab initioinitio accuracy at orders-of-magnitude lower computational expense. Machine-learned interatomic potentials (MLIPs) utilize parametric functionals to represent interatomic interactions. Trained on atomic structure data, MLIPs are developed through a graph neural network (GNN)-based architecture, which directly maps atomic configurations to their corresponding potential energy surface (PES), without the need for handcrafted feature descriptors. Many state-of-the-art (SOTA) GNNs, such as MACE[batatia2022mace], NequIP[batzner20223], eSEN[fu2025learning], and PaiNN[schutt2021equivariant], have been developed to leverage equivariant message passing to achieve high data efficiency. Our agent integrates the open-source Universal Models of Atoms–Medium (UMA-M)[wood2025family], which comprises five specialized models, each uniquely designed for distinct categories of material systems. While the models have exhibited remarkable generalizability across a large chemical space, prediction accuracy for novel compositions may still be constrained by the inherent limitations of the training data, which lack a proper description of charge, spin, and strong $d$-orbital correlations. For specialized systems where these effects are dominant, our agent develops specialized MLIPs, which are typically trained on only a few atom types, but they can be rapidly fine-tuned and uncertainty quantified via ensemble variance methods. Training of MLIPs is implemented in a multi-target learning scheme, where the node features are optimized to simultaneously predict multiple properties. The loss function is given by:
@@
where $\alpha_{E}$, $\alpha_{F}$, and $\alpha_{S}$ are the weight coefficients for total energy, atomic force, and virial stress, respectively, NN denotes the number of training samples, $X_{i}$ and $\hat{X}_{i}$ ($X=E,\mathbf{F},\sigma$) are the true and predicted targets for structure $i$, respectively. Optimized hyperparameters are often determined using Bayesian optimization methods.[martinez2018practical] MLIP ensembles are constructed by training multiple models on identical datasets but with different parameter initializations. The variance in predictions across the ensemble provides a rigorous, $a$ prioripriori metric for assessing model confidence and identifying extrapolation beyond the trained domain[musil2019fast]. Active learning is implemented to ensure transferability of specialized models across the desired chemical space.[podryabinkin2017active, podryabinkin2019accelerating, zhang2019active] The training dataset is iteratively enriched by incorporating new structures that are generated from the conditioned atom types and have the highest epistemic uncertainty.
@@ 8 Physics-Informed Prediction
Many emergent macroscale properties of materials cannot be simply derived from a single DFT geometric optimization calculation, but instead require multiple DFT calculations integrated with physics-informed analysis. Hence, the computational cost can grow exponentially with increasing system size and project scope. In contrast, MLIPs can unlock the unprecedented efficiency in predictive screening while preserving near-abab initioinitio fidelity. To enable higher-level property predictions, our current version provides direct interfaces to Pymatgen[ong2013python], Phonopy, Phono3py[togo2023first, togo2023implementation], and DeePMD-kit.[zeng2023deepmd] Figure 2 contrasts the computational cost scaling between DFT and MLIPs across a range of system sizes for various simulation tasks. These performance benchmarks are averaged from the estimated calculation convergence time observed across a series of Mg-Si-O-H chemical systems. The computational cost of obtaining emergent properties via MLIP inference is 103 $\sim$ 105 times lower than equivalent DFT calculations. It should be noted that the exact magnitude of this advantage can vary with the chemical system, the resolution of the computational methods employed ($e.g.,$ the number of intermediate images in nudged elastic band calculations), and the simulation timescale ($e.g.,$ for molecular dynamics trajectories), yet it consistently reflects a significant reduction in computational expense for MLIP. Crucially, this cost reduction renders many prohibitively expensive calculations ($e.g.,$ phonon vibrations) feasible for high-throughput screening.
@@ 9.1 Problem Framing
Herein, we present a use case of Aethorix v1.0 in resolving a persistent quality control challenge in Portland cement manufacturing. As detailed in Figure 3, the objective of the example use case is to systematically reduce the day-to-day production variability in the compressive strength across both early-age (1-3 day) and long-term (28 day) curing regimes. After consolidating the predominant expert perspectives from a corpus of 18 scholarly works published within the last five years,[fayaz2025industrial, venkatesh2020predictive, pollabauer2024modellbasierte, zanoli2020optimization, tsamatsoulis2024robust, pan2024operation, lyu2024quality, hao4590797control, ali2022machine, teplicka2023evaluation, sutawijaya2021optimizing, farazmand2024effect, saied2020process, costa2020portland, kolesnikov2021review, li2022substance, huilcapi2020analisis, masmali2021implementation] the integrated SciSpace module summarized leading causal factors and proposed established mitigation strategies for raw materials engineering and process control optimization (Figure 3a). Following user interaction, our agent formulated actionable plans for solution development (Figure 3b), prompting the user to select an implementation path based on their prioritized needs. It is noteworthy that the framework supported further interactive dialogue to elaborate on the underlying design principles of each proposed solution.
@@ 9.2 Phase Stabilization through Atomistic Engineering
The elevated calcination temperature of 1730 K necessitated explicit consideration of entropic contributions to phase stability. C3S and C2S are the dominant compositions in cement clinkers. The computed Gibbs free energies of formation for polymorphic C3S and C2S phases at finite temperatures are presented in Figure 4e–g and Figure S1, respectively. The thermodynamic landscape reveals a broad distribution of triclinic and orthorhombic symmetries across both systems. The distribution shifted significantly between room temperature and 1730 K. While experimental identification of exact C3S and C2S crystallographic symmetries remains challenging, our predicted most stable phases exhibit good consistency with the reported industrial clinker polymorphs.[ludwig2015research] At 1730 K, the equilibrium lattice parameters of the most thermodynamically stable C3S and C2S polymorphs are shown as follows:
@@
The early-age properties and long-term performance of concrete are directly governed by the hydration behavior of its clinkers.[claverie2020ab, geng2025investigation] However, the hydration of clinkers involves complex, coupled reaction mechanisms ($e.g.,$ ionic dissolution, degree of hydroxylation, evolution of surface structure, and chemical bond formation),[geng2025investigation, chen2025molecular] making the modeling of the full process at atomistic levels prohibitively expensive. Our agent captured the bulk ionic mobility to approximate hydration reactivity of clinkers. Figure 4f shows the ionic self-diffusivity distribution for the thermodynamically most stable C3S polymorphs, quenched from 1730 K. The self-diffusivity values were derived from short-time molecular dynamics simulations (1 ps) at 300 K via the Einstein relation:
@@
where NN is the number of atoms, $d$ is the dimensionality of diffusion, $\mathbf{r}_{i}$ is the position of atom $i$ at $t$, and the angular brackets denote an ensemble or time average. The self-diffusivity was found to vary by 1–2 orders of magnitude across all thermodynamically most stable polymorphs. Those with low ionic mobility ($e.g.,$ ID 01) were hence identified undesirable and their production was targeted for suppression during synthesis. Figure 4g and S1c show the phase stabilization effect due to presence of low concentration ($\sim$10 mol%) of dopant ions in clinker crystal lattices at 1730 K. The results revealed varying dopant site occupancy preferences across polymorphs, which deviate from empirical trends.[zhang2011development, young1996highly, schwiete1968existenzbereiche] For instance, Na preferentially substitutes the Ca site in the C2S ID10 lattice, whereas it favors the Si site in the C3S ID25 lattice. Most of these substitutions resulted in a reduced free energy at 1730 K, Notable exceptions, such as S or P substituting for Ca, are likely due to the significant charge mismatch introduced by the dopant, which destabilizes the lattice. These are supported by experimental observations.[ludwig2015research] Crucially, K doping provided the strongest stabilization of all dopants for the desired polymorphs, while having a negligible effect on the undesired ones. It was further verified that incorporation of K does not impede ion mobility within the desired phases. Therefore, it was proposed that industrial clinker production strategically maintain K impurity levels at 5–10 mol% to favor the formation of hydration-reactive polymorphs, a direct route to improving the performance of concrete from early ages to long-term service.